# Laser Interactions with Gas Jets: EMP Emission and Nozzle Damage

### v2

@@ III.1 X-ray Photoionisation of the Gas
where the electron density is in cm-3 and the electron temperature is in eV. The radiative contribution of bulk electrons dominates that of hot electrons because $n_{e}T_{e}\gtrsim n_{h}T_{h}$ under the present conditions. For an electron density of $n_{e}=8\times 10^{20}$ cm-3 in fully ionised plasma and a temperature $T_{e}\simeq 250$ keV, we have $P_{\text{Brem}}=6\times 10^{12}$ W cm-3. Accounting for the plasma volume of $V_{p}\sim 6\times 10^{-7}$ cm3 and lifetime $t_{\rm exp}\sim 60$ ps, the total radiated energy is about $\mathcal{E}_{\rm Brem}=P_{\rm Brem}V_{p}t_{\rm exp}\simeq 0.2$ mJ. The spectrum of emission is flat for photons with energy $\varepsilon_{x}\lesssim T_{e}$ and decreases exponentially with a temperature $T_{\rm rad}\simeq T_{e}$ for higher energies.
The mean free path of photons with energies above 3 keV in hydrogen is approximately constant at $\lambda_{x}\rho_{\rm gas}\sim 3$ g cm-2[41]. For a gas jet density of $\rho_{\rm gas}\sim 1.3$ mg cm-3, the absorption length is 20 m and the probability of gas ionisation over a millimetre distance is negligible. The mean free path of lower energy photons strongly decreases as the energy decreases [41], however. It can be approximated by $\lambda_{x}\rho_{\rm gas}\sim 10^{-10}\varepsilon_{x}^{3}$ g cm-2 for hydrogen, where $\varepsilon_{x}$ is the photon energy in eV. The mean free path of photons with the energy of $15-20$ eV - comparable to the hydrogen ionisation potential $U_{\rm ion}=13.6$ eV - is therefore just a few microns.
@@ III.3 Ionisation by Plasma Ions
The stopping power of a representative 0.4 MeV proton in hydrogen is 1200 MeV cm2/g  [45]. The energy needed for the creation of one secondary electron by a fast ion is of order the ionisation potential, $U_{\rm ion}$, and it varies weakly with the ion energy[46, 47]. A 0.4 MeV proton passing through hydrogen gas therefore creates $\sim 6000$ secondary electrons over the 0.5 mm distance between the plasma and the nozzle, losing about 0.1 MeV of its kinetic energy. In a gas volume of $\sim 4\times 10^{-4}$ cm-3, this corresponds to complete ionisation of the ambient gas. This is, however, a relatively long process that takes $\sim 60$ ps.
@@
In conclusion, photoionisation of the gas jet is low and can be neglected. Instead, ionisation by charged particles is produced in three steps: first, at the level of $10^{-4}$ by fast electrons in a few ps, then at the level of $10^{-3}$ by fast ions within 20 ps, then finally by the expanding plasma on a timescale of 60 ps. Free electrons produced during the first two steps provide the seed for a discharge by electrical breakdown, as discussed in the next section.
@@ III.4 Gas Ionisation by Electrical Breakdown
When hot electrons are ejected from the laser plasma, it becomes positively charged with respect to the nozzle. The electric field between the plasma and nozzle may then break down the gas and propagate a discharge. The discharge regime is determined by the product of the gas pressure and electrode separation $p_{\rm gas}h$. Under Vulcan-like conditions, with $p_{\rm gas}h\gtrsim 1$ bar cm, avalanche ionisation will occur if the potential is above the breakdown value of $20-30$ kV  [48]. This is much smaller than the plasma potential of $\sim 25$ MV estimated in Sec. II.
In the avalanche process, the electron density increases exponentially in time from the seed level $n_{e0}/n_{\rm at}\sim 10^{-4}-10^{-3}$ created by fast electrons or ions, $n_{e}(t)=n_{e0}\,\exp(\nu_{\rm ion}t)$, where $\nu_{\rm ion}$ is the characteristic ionization rate. According to Refs. [48, 49], the ionisation rate is $\nu_{\rm ion}\sim 10^{12}$ s-1 for Vulcan-like conditions and the gas can be fully ionised in a few picoseconds. However, the plasma potential decreases with time for two reasons: first, it decreases logarithmically with time because of plasma expansion according to Eq. (5) and second, because the energy deposited during the ionisation process is extracted from the electrostatic energy. The decrease in electrostatic energy is related to the ionisation loss, $\Delta\mathcal{E}\simeq U_{\rm ion}n_{e}V_{\rm gas}$, and the discharge stops when all the available electrostatic energy $\mathcal{E}_{es}$ has been exhausted. This relation defines the maximum ionisation that can be achieved by avalanche breakdown, $n_{e,\max}\sim 4\times 10^{21}$ cm-3, which is larger than the atomic gas density.
In conclusion, avalanche breakdown is initiated from the seed level of ionisation produced by fast electrons or fast ions on a $10-20$ ps timescale. It can fully ionise the gas under the conditions of the Vulcan experiment before the expanding plasma reaches the nozzle.
III.5 Plasma Resistance
@@ IV EMP emission
Here, we estimate the characteristics of EMP emission. The theoretical picture is as follows: the plasma supplies a current limited by the nozzle resistance and the available charge and radiation is emitted from an antenna made out of the conducting parts of the nozzle. Considering the nozzle as a quarter-wavelength dipole, the emission frequency depends on the nozzle length $h_{d}$ as $\nu_{\rm emp}=c/4h_{d}$. For $h_{d}=3$ cm, the emission frequency is 2.5 GHz.
@@
where $Z_{0}=377$ $\Omega$ is the vacuum impedance. For 230 nC accumulated charge and an emission frequency of 2.5 GHz, the relation yields $\sim 5$ mJ of EMP energy and an average current at the antenna frequency[51] of $J_{\text{emp}}=Q\nu_{\rm emp}\sim 580$ A. EMP emission is delayed with respect to the laser: first, because it takes time to ionise the gas and second, because the current needs time to propagate along the antenna.
Knowing the amplitude of the discharge current, we can also estimate the amplitude of the emitted magnetic field at a distance $r$ as $B\simeq\mu_{0}J_{\text{emp}}/2\pi r$. This is equivalent to the maximum B-field (measured over all possible emission angles) in the far-field of a dipole antenna[51]. For a nozzle of height $h_{d}=3$ cm, the peak magnetic field amplitude is $\sim 120$ $\mu$T at a distance of $r=1$ m. The gas discharge is thus expected to produce an electromagnetic pulse with electric field of amplitude $E=cB\approx 36$ kV/m (in the plane wave approximation). This value is higher than the electromagnetic susceptibility threshold for electronics [52].
@@ V Nozzle Damage Mechanism
We consider copper as a representative nozzle material, with a heat capacity $C_{V}=0.39$ J/gK, a melt temperature of 1360 K and a latent heat of melting 0.86 kJ/g. To heat 1 g of Cu to the melting temperature - and to melt it - one needs a total energy of 1.3 kJ. Similar estimates apply to aluminium and iron. The total available electrostatic energy in the plasma of 3 J can melt a maximum of 2 mg Cu. By contrast, the kinetic energy of the expanding plasma is more than 10 times larger. The plasma can therefore cause more damage to the nozzle.
@@ V.2 Heating caused by Plasma Deposition
Based on the estimates of plasma density and volume from Sec. II.1, there are $N_{i}\sim 5\times 10^{14}$ ions at energy $\sim 0.4$ MeV for a combined energy of 40 J. Half of these ions strike the nozzle and deposit an energy of 250 kJ/g. This is more than 100 times higher than the energy required to melt Cu, so plasma heating melts the nozzle easily.
The same conclusion applies to a ceramic nozzle, even with the greater heat resistance. Consider SiC with a melting temperature of 3100 K, 370 J/g latent heat of melting and a specific heat capacity of $0.67-1.4$ J/gK for temperatures between $300-4000$ K [54]. Using these values, we find that $\sim 4$ kJ/g is required to melt SiC. According to the NIST proton stopping catalogue, the stopping power of a 0.5 MeV proton in SiO2 is 290 MeV cm2/g [45]. For a density of 3.2 g cm-3, the rate of energy deposition is 1000 MeV/cm, and the deposition depth is $\sim 5\,\mu$m, for a heated volume of approximately $2\times 10^{-5}$ cm3 and a heated mass of $0.06-0.1$ mg. This is consistent with SRIM Monte Carlo simulations [55] of 0.4 MeV protons incident on SiC, which give a $4\,\mu$m range. 20 J of incident ion energy deposited in a few microns corresponds to heating of $\sim 2$ MJ/g - more than sufficient to melt the ceramic.
@@
We can apply this model of nozzle damage to a Vulcan-style experiment. The amount of energy deposited by the ions is sensitive to the number of ions and their average energy. Taking a vertical gas density profile from measurements by Hicks et al. [26], we estimate the number of ions generated in the laser channel when the laser is focused at different heights above the gas jet. We assume a value of $N_{i}$ at a fixed height of $160\,\mu$m, where the density is maximal, and scale it by the normalised vertical density profile as the laser focus is shifted to different heights above the nozzle. Following the method outlined earlier, we then calculate the ion flux on the nozzle surface and the amount of energy deposited when the laser is focused at different distances from the nozzle. Hicks et al. observed that metallic nozzles were instantly destroyed at a nozzle-focus distance of $400\,\mu$m and virtually undamaged when they were separated by 2 mm. Since the melt threshold for Cu is 1.4 kJ/g, this suggests that the ion energy deposition must reach 2 kJ/g when the laser is focused somewhere between these two distances. We therefore have bounds on the total number of 0.4 MeV plasma ions produced by the laser at a height of $160\,\mu$m, which satisfies $4\times 10^{12}<N_{i}<2\times 10^{14}$. The product of the initial plasma volume and peak gas density gives a maximum electron number of $\sim 5\times 10^{14}$, which may be considered an upper limit on the number of accelerated ions.
Figure 3: Ion energy deposited per mass of nozzle material (Cu or Fe) as a function of laser-nozzle separation for the number of ions $N_{i}=4\times 10^{12}$ (a) and $2\times 10^{14}$ (b) normalised to the plasma height of $160\,\mu$m. The red dot represents the theoretical melt threshold for a Cu nozzle. Red and green vertical dashed lines represent observed distances where a steel nozzle was destroyed and survived, respectively. The ion energy is 0.5 MeV.
The situation is illustrated graphically in Fig. 3. The left panel shows the ion energy deposited in the nozzle as a function of distance, assuming $4\times 10^{12}$ plasma ions at 0.5 MeV when the laser is focused at a height of $160\,\mu$m. The right panel shows the energy deposition for $2\times 10^{14}$ plasma ions at a height of $160\,\mu$m above the nozzle. The amount of nozzle heating increases as the laser approaches the peak gas density at $\sim 160\,\mu$m and then drops steeply as the laser is focused further away from the nozzle. This is caused partially by the reduced solid angle subtended by the nozzle and also the changing gas density profile. Accounting for ion collisions in the gas before they reach the nozzle means that less overall energy is deposited in the nozzle, but the energy per unit mass of heated material increases due to a higher ion stopping power at lower energies.
@@ VI.1 Ion Acceleration Experiment on VEGA-3
An experiment was conducted on the VEGA-3 laser system with the aim of using PW laser pulses to trigger laser channelling in inert gases (He, N2, He-N2) and thereby learn more about ion acceleration mechanisms [24, 25, 25].
@@
To measure the electron density in the plasma channel, a folding-wave interferometer was used to take optical interferograms of the ionised gas at different temporal delays relative to the drive laser [57]. Figure 4b shows a He density distribution from the experiment taken $\simeq 150$ ps after the arrival of the laser, revealing a laser channel with average radius $\simeq 70\leavevmode\nobreak\ \mu$m. Modelling the expansion as a Sedov-Taylor cylindrical blast wave [58, 59], $R(t)=\xi_{0}(\mathcal{E}_{b}t^{2}/\rho_{\rm gas}l_{p})^{1/4}$ gives an energy of $\mathcal{E}_{b}\simeq 1.3$ J (taking $\xi_{0}=0.9$ [59]) deposited in the channel over its length of $l_{p}\simeq 0.5$ mm, which is less than 20% the deposited laser energy of 7.3 J. Such a significant difference is explained by the fact that the expanding plasma is essentially collisionless. Plasma ions propagate through the ambient gas and ionise atoms without creating a density compression - only a small portion of the slow ion population contribute to shock wave formation.
2D PIC simulations in CALDER[25, 29] and experimental interferograms of the gas at the point of laser arrival suggest that, by the time the laser has travelled through the low-density wings of the gas density profile and arrived in the central high-density region, the laser has become significantly defocused. For a gas density of 4.5 bar and a plasma channel radius of $20\leavevmode\nobreak\ \mu$m, the plasma volume is $\sim 6\times 10^{-7}$ cm3, the laser energy deposited in the gas is 7.3 J, the bulk electron temperature is $T_{e}\simeq 140$ keV and the pressure is $\sim$50 Mbar (see Tab. 1).
The stopping power of 140 keV electrons in He is $\sim 3$ MeV cm2/g [42], which corresponds to a mean free path $>50$ cm for our gas density. The Debye radius, however, is of order $0.1\leavevmode\nobreak\ \mu$m, so bulk electrons remain in the plasma due to the Coulomb attraction and the plasma expansion is quasi-neutral. Bremsstrahlung losses are negligible on this timescale, so the plasma expands adiabatically.
@@
Applying the analysis of gas ionisation described in Sec. II, we find that the photoionisation probability is very low - about $1.6\times 10^{-7}$ the level of ionisation by fast electrons is two orders of magnitude larger at $\sim 3.3\times 10^{-5}$ and the dominant ionisation processes are electrical breakdown and ions in the expanding plasma. The plasma ionisation time is therefore within a few tens of ps. The plasma discharge results in EMP emission with energy $\sim$0.17 mJ, which corresponds to an electric field of $\sim 6.3$ kV/m and a magnetic field of $21\,\mu$T at 1 m from the source. Nozzle damage is due to energy deposition of a few joules by the plasma ions.
Figure 5: Interferograms of the VEGA-3 laser interacting with an N2 gas ejected from a short-focus shock nozzle at different probe times relative to the arrival of the pump beam: (a) Probe time $t\sim 40$ ps (b) Probe time $t\sim 90$ ps.
Progress of gas ionisation between the laser and nozzle was measured with a $2\omega$ pick-off beam directed transversely to the laser channel for interferometric imaging. Figure 5 shows interferograms taken at two different times after the arrival of the laser pulse. The gas is N2, here. Figure 5(a) corresponds to the earliest probing time measured after the laser pulse arrival at $t=0$ ps. The shadow of the gas is visible in both images, suggesting that the entire gas volume can be ionised within 40 ps. The speed of the ionisation front therefore exceeds $\sim 20\,\mu$m/ps, which is consistent with gas ionisation by electrical breakdown.
To monitor ion emission during the experiment, five diamond time-of-flight detectors [60] were placed in multiple different locations around the vacuum chamber ($0^{\circ}$, ±$\pm 17^{\circ}$, $4.85^{\circ}$, $5.5^{\circ}$ with respect to the laser axis, always inclined at $9^{\circ}$ to the horizontal plane of the laser) and a Thomson parabola spectrometer was placed along the laser axis in the same horizontal plane as the laser. When the diamond detectors were placed at ±$\pm 17^{\circ}$, nitrogen ions with energy up to $\sim 80$ MeV were measured. For shots on He gas, $\alpha$-particles were accelerated to 0.7 MeV/u with $10^{9}$ particles per steradian, while shots on N2 gas produced ions up to 5.7 MeV/u with $10^{7}$ particles per steradian. These ions were measured in the laser forward direction (within a ±$\pm 17^{\circ}$ cone)[25].

### v3

@@ III.1 X-ray Photoionisation of the Gas
where the electron density is in cm-3 and the electron temperature is in eV. The radiative contribution of bulk electrons dominates that of hot electrons because $n_{e}T_{e}\gtrsim n_{h}T_{h}$ under the present conditions. For an electron density of $n_{e}=8\times 10^{20}$ cm-3 in fully ionised plasma and a temperature $T_{e}\simeq 250$ keV, we have $P_{\text{Brem}}=6\times 10^{12}$ W cm-3. Accounting for the plasma volume of $V_{p}\sim 6\times 10^{-7}$ cm3 and lifetime $t_{\rm exp}\sim 60$ ps (see Sec. II.3), the total radiated energy is about $\mathcal{E}_{\rm Brem}=P_{\rm Brem}V_{p}t_{\rm exp}\simeq 0.2$ mJ. The spectrum of emission is flat for photons with energy $\varepsilon_{x}\lesssim T_{e}$ and decreases exponentially with a temperature $T_{\rm rad}\simeq T_{e}$ for higher energies.
The mean free path of photons with energies above 3 keV in hydrogen is approximately constant at $\lambda_{x}\rho_{\rm gas}\sim 3$ g cm-2[49]. For a gas jet density of $\rho_{\rm gas}\sim 1.3$ mg cm-3, the absorption length is 20 m and the probability of gas ionisation over a millimetre distance is negligible. The mean free path of lower energy photons strongly decreases as the energy decreases [49]. It can be approximated as $\lambda_{x}\rho_{\rm gas}\sim 10^{-10}\varepsilon_{x}^{3}$ g cm-2 for hydrogen, where $\varepsilon_{x}$ is the photon energy in eV. The mean free path of photons with the energy of $15-20$ eV - comparable to the hydrogen ionisation potential $U_{\rm ion}=13.6$ eV - is therefore just a few microns.
@@ III.3 Ionisation by Plasma Ions
The stopping power of a representative 0.4 MeV proton in hydrogen is 1200 MeV cm2/g  [53]. The energy needed for the creation of one secondary electron by a fast ion is of order the ionisation potential, $U_{\rm ion}$, and it varies weakly with the ion energy [54, 55]. A 0.4 MeV proton passing through hydrogen gas, therefore, creates $\sim 6000$ secondary electrons over the 0.5 mm distance between the plasma and the nozzle, losing about 0.1 MeV of its kinetic energy. In a gas volume of $\sim 4\times 10^{-4}$ cm-3, this corresponds to complete ionisation of the ambient gas. This is, however, a relatively long process that takes $\sim 60$ ps.
@@ III.4 Gas Ionisation by Electrical Breakdown
When hot electrons are ejected from the laser plasma, it becomes positively charged with respect to the nozzle. The electric field between the plasma and nozzle may then break down the gas and propagate a discharge. The discharge regime is determined by the product of the gas pressure and electrode separation $p_{\rm gas}h$. Under Vulcan-TAP conditions, with $p_{\rm gas}h\gtrsim 1$ bar cm, avalanche ionisation will occur if the potential is above the breakdown value of $20-30$ kV  [56]. This is much smaller than the plasma potential of $\sim 25$ MV estimated in Sec. II.
In the avalanche process, the electron density increases exponentially in time from the seed level $n_{e0}/n_{\rm at}\sim 10^{-4}-10^{-3}$ created by fast electrons or ions, $n_{e}(t)=n_{e0}\,\exp(\nu_{\rm ion}t)$, where $\nu_{\rm ion}$ is the characteristic ionization rate. According to Refs. [56, 57], the ionisation rate is $\nu_{\rm ion}\sim 10^{12}$ s-1 for Vulcan conditions and the gas can be fully ionised in a few picoseconds. However, the plasma potential decreases with time for two reasons: first, it decreases logarithmically with time because of plasma expansion according to Eq. (5) and second, because the energy deposited during the ionisation process is extracted from the electrostatic energy. The decrease in electrostatic energy is related to the ionisation loss, $\Delta\mathcal{E}\simeq U_{\rm ion}n_{e}V_{\rm gas}$, and the discharge stops when all the available electrostatic energy $\mathcal{E}_{es}$ has been exhausted. This relation defines the maximum ionisation that can be achieved by avalanche breakdown, $n_{e,\max}\sim 4\times 10^{21}$ cm-3, which is larger than the atomic gas density.
III.5 Conclusions on Ionization Processes and Plasma Resistance
In conclusion, photoionisation of the gas jet is low and can be neglected. Instead, ionisation by charged particles is produced in three steps: first, at the level of $10^{-4}$ by fast electrons in a few ps, then at the level of $10^{-3}$ by fast ions within 20 ps, then finally by the expanding plasma on a timescale of 60 ps. Free electrons produced during the first two steps provide the seed for a discharge by electrical breakdown. Avalanche breakdown is initiated from the seed level of ionisation produced by fast electrons or fast ions on a $10-20$ ps timescale. It can fully ionise the gas under the conditions of the Vulcan experiment before the expanding plasma reaches the nozzle.
@@ IV EMP emission
Here, we estimate the characteristics of EMP emission. The theoretical picture is as follows: the plasma supplies a current limited by the nozzle resistance and the available charge, and radiation is emitted from an antenna made out of the conducting parts of the nozzle. Considering the nozzle as a quarter-wavelength dipole, the emission frequency depends on the nozzle length $h_{d}$ as $\nu_{\rm emp}=c/4h_{d}$. For $h_{d}=3$ cm, the emission frequency is 2.5 GHz.
@@
where $Z_{0}=377$ $\Omega$ is the vacuum impedance. For 230 nC accumulated charge and an emission frequency of 2.5 GHz, the relation yields $\sim 5$ mJ of EMP energy and an average current at the antenna frequency [27] of $J_{\text{emp}}=Q\nu_{\rm emp}\sim 580$ A. EMP emission is delayed with respect to the laser: first, because it takes time to ionise the gas and second because the current needs time to propagate along the antenna.
Knowing the amplitude of the discharge current, we can also estimate the amplitude of the emitted magnetic field at a distance $r$ as $B\simeq\mu_{0}J_{\text{emp}}/2\pi r$, where $\mu_{0}$ is the permeability of free space. This is equivalent to the maximum magnetic field (measured over all possible emission angles) in the far-field of a dipole antenna [27]. For a nozzle of height $h_{d}=3$ cm, the peak magnetic field amplitude is $\sim 120$ $\mu$T at a distance of $r=1$ m. The gas discharge is thus expected to produce an electromagnetic pulse with electric field of amplitude $E=cB\approx 36$ kV/m (in the plane wave approximation). This value is higher than the electromagnetic susceptibility threshold for electronics [59].
@@ V Nozzle Damage Mechanism
We consider copper as a representative nozzle material, with a heat capacity $C_{V}=0.39$ J/gK, a melt temperature of 1360 K and a latent heat of melting 0.86 kJ/g. To heat 1 g of Cu to the melting temperature - and to melt it - one needs a total energy of 1.3 kJ. Similar estimates apply to aluminium and iron. The total available electrostatic energy in the plasma of 3 J can melt a maximum of 2 mg Cu. By contrast, the total kinetic energy of the expanding plasma ions is more than 10 times larger. The plasma can therefore cause more damage to the nozzle.
@@ V.2 Heating caused by Plasma Deposition
Based on the estimates of plasma density and volume from Sec. II.3, there are $N_{i}\sim 5\times 10^{14}$ ions at energy $\sim 0.4$ MeV for a combined energy of 40 J. Half of these ions strike the nozzle and deposit an energy of 250 kJ/g. This is more than 100 times higher than the energy required to melt Cu, so plasma heating melts the nozzle easily.
The same conclusion applies to a ceramic nozzle, even with the greater heat resistance. Consider SiC with a melting temperature of 3100 K, 370 J/g latent heat of melting and a specific heat capacity of $0.67-1.4$ J/gK for temperatures between $300-4000$ K [61]. Using these values, we find that $\sim 4$ kJ/g is required to melt SiC. According to the NIST proton stopping catalogue, the stopping power of a 0.5 MeV proton in SiO2 is 290 MeV cm2/g [53]. For a density of 3.2 g cm-3, the rate of energy deposition is 1000 MeV/cm, and the deposition depth is $\sim 5\,\mu$m, for a heated volume of approximately $2\times 10^{-5}$ cm3 and a heated mass of $0.06-0.1$ mg. This is consistent with SRIM Monte Carlo simulations [62] of 0.4 MeV protons incident on SiC, which give a $4\,\mu$m range. 20 J of incident ion energy deposited in a few microns corresponds to heating of $\sim 200$ kJ/g - more than sufficient to melt the ceramic.
@@
We apply this model of nozzle damage to a Vulcan-TAP experiment. The amount of energy deposited by the ions is sensitive to the number of ions and their average energy. Taking a vertical gas density profile from measurements by Hicks et al. [26], we estimate the number of ions generated in the laser channel when the laser is focused at different heights above the gas jet. We assume a value of $N_{i}$ at a fixed height of $160\,\mu$m, where the density is maximal, and scale it by the normalised vertical density profile as the laser focus is shifted to different heights above the nozzle. Following the method outlined earlier, we then calculate the ion flux on the nozzle surface and the amount of energy deposited when the laser is focused at different distances from the nozzle. In the experiment, metallic nozzles were instantly destroyed at a nozzle-focus distance of $400\,\mu$m and virtually undamaged when they were separated by 2 mm. Since the melt threshold for Cu is 1.4 kJ/g, this suggests that the ion energy deposition must reach $\sim 2$ kJ/g when the laser is focused somewhere between these two distances. These observations give bounds on the total number of 0.5 MeV plasma ions produced by the laser at a height of $160\,\mu$m, which satisfies $4\times 10^{12}<N_{i}<2\times 10^{14}$. The product of the initial plasma volume and peak gas density gives a maximum electron number of $\sim 5\times 10^{14}$, which may be considered an upper limit on the number of accelerated ions.
Figure 4: Ion energy deposited per mass of nozzle material as a function of laser-nozzle separation for the number of ions $N_{i}=4\times 10^{12}$ (a) and $2\times 10^{14}$ (b) normalised to the plasma height of $160\,\mu$m. The red dot represents the theoretical melt threshold for a Cu nozzle. Red and green vertical dashed lines represent observed distances where a steel nozzle was destroyed and survived, respectively. The ion energy is 0.5 MeV.
The situation is illustrated graphically in Fig. 4. Panel (a) shows the ion energy deposited in the nozzle as a function of distance, assuming $4\times 10^{12}$ plasma ions at 0.5 MeV when the laser is focused at a height of $160\,\mu$m. Panel (b) shows the energy deposition for $2\times 10^{14}$ plasma ions at a height of $160\,\mu$m above the nozzle. The amount of nozzle heating increases as the laser approaches the peak gas density at $\sim 160\,\mu$m and then drops steeply as the laser is focused further away from the nozzle. This is caused partially by the reduced solid angle subtended by the nozzle and also the changing gas density profile. Accounting for ion collisions in the gas before they reach the nozzle means that less overall energy is deposited in the nozzle, but the energy per unit mass of heated material increases due to a higher ion stopping power at lower energies.
@@ VI.1 Ion Acceleration Experiment on VEGA-3
An experiment was conducted on the VEGA-3 laser system with the aim of using petawatt laser pulses to trigger laser channelling in inert gases (He, N2, He-N2) and thereby learn more about ion acceleration mechanisms [25, 24].
Figure 5: Schematic diagrams of two experiments conducted on the VEGA-3 laser system. (a) Set-up described in Sec. VI.1. A Thomson parabola spectrometer (TP), two electron spectrometers (E-Spec) and four diamond time-of-flight detectors (Diamond ToF) are represented by coloured boxes. (b) Set-up described in Sec. VI.2. Three Thomson parabola spectrometers are placed at $0^{\circ}$, $60^{\circ}$ and $90^{\circ}$ to the laser axis. The B-dot probe was positioned at variable distances ($\sim 2-3$ m) from the gas target, with its measurement axis horizontal and orthogonal to the line-of-sight axis. A camera was used to take images of the gas at twice the laser fundamental frequency. Further details of these experiments can be found in Refs. [25, 23].
@@
To measure the electron density in the plasma channel, a folding-wave interferometer was used to take optical interferograms of the ionised gas at different temporal delays relative to the drive laser [64]. Figure 6b shows a He density distribution from the experiment taken $\simeq 150$ ps after the arrival of the laser, revealing a laser channel with average radius $\simeq 70\leavevmode\nobreak\ \mu$m. Interferograms could not be recovered at earlier probing times because the channel boundaries were obscured by plasma self-emission. Modelling the expansion as a Sedov-Taylor cylindrical blast wave [65, 66], $R(t)=\xi_{0}(\mathcal{E}_{b}t^{2}/\rho_{\rm gas}l_{p})^{1/4}$, gives an energy of $\mathcal{E}_{b}\simeq 1.3$ J (taking $\xi_{0}=0.9$ [66]) deposited in the channel over its length of $l_{p}\simeq 0.5$ mm, which is less than 20% the deposited laser energy of 7.3 J. Such a significant difference is explained by the fact that the expanding plasma is essentially collisionless. Plasma ions propagate through the ambient gas and ionise atoms without creating a density compression - only a small portion of the slow ion population contributes to shock wave formation.
2D PIC simulations in CALDER [25, 28] and experimental interferograms of the gas at the point of laser arrival suggest that, by the time the laser has travelled through the low-density wings of the gas density profile and arrived in the central high-density region, the laser has become significantly defocused. For a gas density of 4.5 bar and a plasma channel radius of $20\leavevmode\nobreak\ \mu$m, the plasma volume is $\sim 6\times 10^{-7}$ cm3, the laser energy deposited in the gas is 7.3 J, the bulk electron temperature is $T_{e}\simeq 140$ keV and the pressure is $\sim 50$ Mbar, see Tab. 3).
The stopping power of 140 keV electrons in He is $\sim 3$ MeV cm2/g [50], which corresponds to a mean free path of more than 50 cm for our gas density. The Debye radius, however, is of order $0.1\leavevmode\nobreak\ \mu$m, so bulk electrons remain in the plasma due to the Coulomb attraction and the plasma expansion is quasi-neutral. Bremsstrahlung losses are negligible on this timescale, so the plasma expands adiabatically.
@@
Applying the analysis of gas ionisation described in Sec. II, we find that the photoionisation probability is very low - about $1.6\times 10^{-7}$ the level of ionisation by fast electrons is two orders of magnitude larger at $\sim 3.3\times 10^{-5}$ and the dominant ionisation processes are electrical breakdown and ions in the expanding plasma. The plasma ionisation time is, therefore, within a few tens of ps. The plasma discharge results in EMP emission with energy $\sim 0.17$ mJ, which corresponds to an electric field of $\sim 6.3$ kV/m and a magnetic field of $21\,\mu$T at 1 m from the source. Nozzle damage is due to the energy deposition of a few joules by the plasma ions.
Figure 7: Interferograms of the VEGA-3 laser interacting with an N2 gas ejected from a short-focus shock nozzle. Probe times relative to the arrival of the pump beam are 40 ps (a) and 90 ps (b). The laser intensity is $10^{20}$ W/cm2 and the gas density is $10^{20}$ cm-3.
Progress of gas ionisation between the laser and nozzle was measured with a pick-off beam converted in the second harmonic and directed transversely to the laser channel for interferometric imaging. Figure 7 shows interferograms taken at two different times after the arrival of the laser pulse. Panel (a) corresponds to the earliest probing time measured after the laser pulse arrival. The shadow of the gas is visible in both images, suggesting that the entire gas volume can be ionised within 40 ps. The speed of the ionisation front, therefore, exceeds $\sim 20\,\mu$m/ps, which is consistent with gas ionisation by electrical breakdown.
To monitor ion emission during the experiment, five diamond time-of-flight detectors [67] were placed in multiple different locations around the vacuum chamber ($0^{\circ}$, ±$\pm 17^{\circ}$, $4.85^{\circ}$, $5.5^{\circ}$ with respect to the laser axis, always inclined at $9^{\circ}$ to the horizontal plane of the laser) and a Thomson parabola spectrometer was placed along the laser axis in the same horizontal plane as the laser. When the diamond detectors were placed at ±$\pm 17^{\circ}$, nitrogen ions with energy up to $\sim 80$ MeV were measured. For shots on He gas, $\alpha$-particles were accelerated to 0.7 MeV/u with $10^{9}$ particles per steradian, while shots on N2 gas produced ions up to 5.7 MeV/u with $10^{7}$ particles per steradian. These ions were measured in the laser forward direction within a ±$\pm 17^{\circ}$ cone [25]. Sample He spectra from the time-of-flight diagnostics can be found in Fig. 3b. We expect higher yields perpendicular to the laser axis.